OpenStudy (anonymous):

What would this be in an exponential equation? In(x+y)=5

OpenStudy (anonymous):

exponentials have exponents. if you convert this in terms of e than it will be.

OpenStudy (debbieg):

\(\Large \log_{b}x=y \) means that \(\Large x=b^y \) The output of the log is the exponent of the exponential form; the base of the log is the base; and the input of the log is the output of the exponential form. Just re-write as an exponential, putting all the pieces in the appropriate place! :)

OpenStudy (anonymous):

is In incorporated with the exponential equation form? I'm having a bit of trouble, Inlog5x=y? log5x=y? @DebbieG

OpenStudy (anonymous):

you dont have to use log. just change the function you have right now to e^5=(x+y) e^5-x=y

OpenStudy (anonymous):

log is base 10 ln is base e.

OpenStudy (anonymous):

log base e is ln

OpenStudy (anonymous):

Thank you so much!